OpenStudy (anonymous):

What is the center and radius of the circle with equation (x - 5)2 + (y + 3)2 = 16?

OpenStudy (amistre64):

given a center (a,b) a circle has the form: (x-a)^2 + (y-b)^2 = r^2

OpenStudy (amistre64):

-a = -5 -b = 3 r^2 = 16 the rest is elementary

OpenStudy (anonymous):

Okasy Thacks I Get center (5, -3); radius = 4?